1)        Johnson Jets is considering two mutually exclusive projects. Project A has an up-front cost...

1)

       Johnson Jets is considering two mutually exclusive projects. Project A has an up-front cost of $122,000 (CF0 = -122,000), and produces positive after-tax cash inflows of $30,000 a year at the end of each of the next six years. Project B has an up-front cost of $60,000(CF0 = -60,000) and produces after-tax cash inflows of $20,000 a year at the end of the next four years. Assuming the cost of capital is 10.5%,
1. Compute the equivalent annual annuity of project A in box 1. Round the EAA to a whole dollar without the dollar sign or comma, e.g., 3452 (non-negative number)
2. Compute the equivalent annual annity of project B in box 2. The same format as box 1.
3. Decide which project to undertake in box 3, either Project A or Project B.

Question 17 options:

Blank # 1

Blank # 2

Blank # 3

2)

There are two alternative machines for a manufacturing process. Both machines have the same output rate, but they differ in costs. Machine A costs $20,000 to set up and $8,000 per year to operate. It must be completely replaced every 3 years, and it has no salvage value. Machine B costs $50,000 to set up and $2,160 per year to operate. It should last for 5 years and has no salvage value. The costs of two machines are shown below.

0

1

2

3

4

5

Machine A

20,000

8,000

8,000

8,000

Machine B

50,000

2,160

2,160

2,160

2,160

2,160


Assuming the cost of capital is 10%,
1. find the equivalent annual cost of Machine A in Box 1. Round it to a whole dollar, and no comma or the dollar sign.
2. find the EAC of Machine B in Box 2. The same format as box 1.
3. Based on the equivalent annual cost method, type in Box 3 which machine do you recommend, Machine A or Machine B.

Question 18 options:

Blank # 1

Blank # 2

Blank # 3
Answer #1

First Question:

the equivalent annual annuity = Annual Cash inflow - CF0 x i / [1 - (1 + i)-n]

Part (1) the equivalent annual annuity for project A = Annual Cash inflow - CF0 x i / [1 - (1 + i)-n]

= 30,000 - 122,000 x 10.5% / [1 - (1 + 10.5%)-6] = 1576 (Please feed this in blank 1)

Part (2) the equivalent annual annuity for project B = Annual Cash inflow - CF0 x i / [1 - (1 + i)-n]

= 20,000 - 60,000 x 10.5% / [1 - (1 + 10.5%)-4] = 866 (Please feed this in blank 2)

Part (3) Since equivalent annual annuity of A > that of B, choose A (Please feed this in blank 3)

Second Question

the equivalent annual cost = Annual outflow + CF0 x i / [1 - (1 + i)-n]

Part (1) the equivalent annual cost for machine A = Annual outflow + CF0 x i / [1 - (1 + i)-n]

= 8,000 + 20,000 x 10% / [1 - (1 + 10%)-3] = 16042 (Please feed this in blank 1)

Part (2) the equivalent annual cost for machine B = Annual outflow + CF0 x i / [1 - (1 + i)-n]

= 2,160 + 50,000 x 10% / [1 - (1 + 10%)-3] = 22266 (Please feed this in blank 2)

Part (3) Since equivalent annual cost of A < that of B, choose A (Please feed this in blank 3)